# Do LVLMs Understand Charts? Analyzing and Correcting Factual Errors in Chart Captioning

### v1

@@ 4 Methodology
To address the factual errors identified in generated captions, we propose a framework, C2TFec, that capitalizes a two-step correction process that offers interpretability. C2TFec first transforms input charts into structured data tables (Section 4.1) and then refines captions by rectifying any discrepancies using the tabular data (4.2). Our deep dive into the performance of a joint framework, “DePlot + GPT-4”, revealed that a notable proportion of errors in caption generation originated from the DePlot component. To mitigate this, we have developed a superior chart-to-table conversion model that builds upon a stronger base model, UniChart, vastly fortified with expansive fine-tuning datasets. The primary advantage of this refined methodology is its ability to harness the strong reasoning strengths of GPT-4 to faithfully correct errors, thereby significantly upscaling the factuality of the generated captions. [^6] Here, we do not consider approaches based on LVLMs due to their tendency towards factual errors as evidenced in Figure 1.
@@
Table 4: Correction performance of different models on the Chocolate dataset. ChartVE measures factuality by computing the entailment probability from each chart to the corresponding caption sentences. GPT-4V and Bard, when used as evaluation metrics, rate each chart caption as factually consistent with the chart or not. Levenshtein computes the edit distance between the corrected caption and the original caption (denoted as “N/A”). Metrics scores are shown separately for each of the three data splits based on captioning model source. The highest and second highest performing models per evaluation metric and split are highlighted in boldface and underlines respectively.
@@ 4.1 Chart-To-Table Conversion
The training data for our chart-to-table model is sourced from datasets including VisText Tang et al. (2023a), Chart-to-text Kantharaj et al. (2022), ChartQA Masry et al. (2022), and PlotQA Methani et al. (2020b), where we repurpose original charts and underlying data tables for our model’s training. We collected a total of 65K instances with a train:dev:test split of 61:2:2. In alignment with DePlot Liu et al. (2023a), our model is also trained to generate chart titles, enhancing its ability to contextualize the data represented in table form. Let $\mathcal{M}$ denote our proposed model. For a given chart figure $\mathcal{E}$, the model autoregressively generates a chart title $\mathcal{T}$ and a corresponding table $\mathcal{A}$ (i.e. $\mathcal{T},\mathcal{A}=\mathcal{M}(\mathcal{E})$).
@@ 4.2 Table-based Error Rectification
With the input chart now converted into structured tabular data, the second phase targets identifying and fixing factual inconsistencies in the initial caption $\mathcal{C}$. Our approach exploits the reasoning capacity of large language models, using GPT-4 to logically analyze the alignment between $\mathcal{C}$ and the constructed table $\mathcal{A}$. Specifically, GPT-4 first provides an explanatory breakdown of detected factual errors in $\mathcal{C}$ based on the table contents. It then uses this explanation to produce an edited caption $\mathcal{\hat{C}}$ remedying the identified inconsistencies. This transparent process enables users to validate the reasoning behind each correction.
Overall, the proposed C2TFec framework aims to produce verified chart captions by systematically isolating the language generation component from the analytical factual verification, taking advantage of the complementary strengths of separate vision and language models tailored to their respective domains. We hypothesize that the intermediate symbolic representation can act as an effective bridge to enforce and validate factual consistency.
@@ 5 Experimental Settings
To evaluate our proposed C2TFec for factual error correction in chart captions, we conduct experiments on our annotated Chocolate dataset.
@@
We utilize the 1,187 chart-caption pairs from the Chocolate dataset, comprising charts paired with both factually consistent and inconsistent captions generated by various models. The captions contain a diverse range of factual errors annotated by human judges, allowing comprehensive assessment. As described in Section 2, the dataset is divided into Lvlm, Llm, and Ft splits based on the caption models.
@@
Since Chocolate does not comprise of training data, we compare C2TFec against competitive baselines that are capable of performing strong zero-shot reasoning. This includes state-of-the-art open-source and proprietary LVLMs, LLaVA-1.5-13B Liu et al. (2023c), GPT-4V OpenAI (2023b), and Bard Google (2023a), as well as a LLM-based pipeline, DePlot + GPT-4 Liu et al. (2023a); OpenAI (2023a). For a fairer comparison between our approach and DePlot, we continue fine-tuning DePlot for an additional 5,000 steps on VisText, an approach which has been shown effective for adapting models to unseen domains Huang et al. (2023b). We denote this model as $\text{DePlot}_{\text{CFT}}$. The prompts used for each model are described in Appendix E.
@@
We assess the factual consistency between corrected captions and input charts using the ChartVE metric, as well as GPT-4V and Bard, according to our recommendations in Section 3.2. In addition, since we require corrections to be made with as few edits as possible, we measure the number of edits using the Levenshtein distance Levenshtein et al. (1966).
@@ 6.1 Main Results
The main results presented in Table 4 demonstrate that our proposed C2TFec framework achieves the best performance in terms of factual consistency on two out of three splits of the Chocolate dataset. Specifically, Chocolate attains the highest scores on the Lvlm and Ft splits, and achieves the second-best performance on the LLM. This indicates that the two-step process of first transforming charts into structured data tables and then rectifying factual inconsistencies using table-caption alignment is an effective strategy.
An interesting observation is that Bard struggles on the Lvlm split. Combined with the negative correlation with human judgments of factuality that we observed in Table 3, we can conclude that Bard is unreliable in identifying factual errors in chart captions. Additionally, when used as an evaluator, GPT-4V tends to assign unduly positive factuality scores to its own corrected outputs. On the Lvlm and Ft splits, GPT-4V rates its own performance as significantly higher than all other models (see Table 6), while other metrics show GPT-4V lagging behind C2TFec. This suggests GPT-4V may suffer from the self-enhancement bias Zheng et al. (2023), failing to reliably identify factual errors in its own generations when served as an evaluation metric. Hence, to further verify the effectiveness of our proposed approach, we conduct human evaluations in Section 6.2.
@@ 6.2 Human Evaluation
Our human evaluations aim to compare the performance of C2TFec and GPT-4V. The annotation tasks for human evaluation are the same as those presented to the annotators in Section 2: identifying various types of factual errors. Hence, the evaluation are conducted by the same annotators that perform the error analysis in Section 2. We sampled 30 charts from each split of Lvlm, Llm, and Ft. For each chart, human judges are presented with a caption generated by one of the models.
Figure 3 encapsulates the conclusive edge of C2TFec over GPT-4V across every error category. Notably, C2TFec demonstrates a marked reduction in Value Errors, surpassing GPT-4V by over a 20% margin within the Lvlm and Llm splits, and also half the overall error rate. Equally impressive was the near elimination of Trend Errors post-correction with C2TFec This underscores C2TFecs robust capability for accurate error rectification pertaining to reasoning regarding axes, including Label, Value, and Trend errors. A qualitative example is shown in Figure 4 to provide intuition on the superior correction ability of C2TFec over GPT-4V. We have also observed that the reason GPT-4V cannot correct these types of errors well is due to its inability to deduce the values of the data points from the input chart (see Figure 7).
In contrast, GPT-4V exhibited a comparative proficiency in addressing Out-of-context Errors, involving information extraneous to the chart’s scope. However, GPT-4V seemed challenged in rectifying errors within captions generated by itself, particularly within the Lvlm split. This observation echoes recent findings on LLMs’ inability to self-correct Huang et al. (2023a); Valmeekam et al. (2023), we find that LVLMs also cannot perform self-correction. More importantly, our human evaluation results, combined with our findings in Table 4 and Table 6, reflect that GPT-4V is subject to serious self-enhancement bias. Consequently, although GPT-4V’s capabilities are formidable, we recommend not using them to assess their own outputs.
@@ 7.1 Chart Captioning
Chart captioning is essential for accurately interpreting and communicating the information conveyed by chart images, particularly in news articles and social media, where factuality is imperative to prevent misinformation. While current datasets like FigureQA Kahou et al. (2017), DVQA Kafle et al. (2018), PlotQA Methani et al. (2020c), VisText Tang et al. (2023a), and Chart-to-Text Kantharaj et al. (2022) offer chart image descriptions and question-answer pairs to train models, advancements in vision-language models like ChartT5 Zhou et al. (2023), MatCha Liu et al. (2023b), and UniChart Masry et al. (2023) have largely prioritized relevance and fluency over factual accuracy. Our work provides a rigorous characterization of factual errors in chart captioning and comparisons of methods to address this gap. By focusing on faithfulness and correction, we complement the emphasis of prior work and aim to produce more trustworthy chart captions.
@@ 7.2 Factual Error Correction
Previous research on based factual error correction has focused on two textual tasks: text summarization and fact-checking. Within text summarization, the bulk of work, such as Dong et al. (2020), has been carried out in the news domain and often involves methods that substitute inconsistent entities from the source text. Some studies have enhanced this approach through entity-replacement reranking techniques Chen et al. (2021), autoregressive models for rewriting and perturbation filtering Cao et al. (2020); Zhu et al. (2021); Adams et al. (2022), and editing strategies that focus on selective deletion Wan and Bansal (2022). In contrast, Fabbri et al. (2022) employed sentence compression datasets to train their models. More recently, Gao et al. (2023) have expanded the focus of these studies to include dialogue summarization.
Moving to the domain of fact-checking, this area has experienced a flurry of activity, particularly with the increased attention on combating misinformation Fung et al. (2021); Wu et al. (2022); Fung et al. (2022); Huang et al. (2023d, c); Qiu et al. (2023). Early approaches train a distantly supervised model that involves a masker and a corrector Shah et al. (2020); Thorne and Vlachos (2021). Thorne and Vlachos (2021) made significant strides by developing the first factual error correction dataset for fact-checking, thus enabling fully supervised training for error correctors. Recently, Huang et al. (2023b) propose an interpretable framework that breaks down the process of fact-checking into individual components. Our study builds on these insights and extends them to a multimodal context, which challenges models to understand the chart images and the consistency between different modalities.
@@ 8 Conclusion
Our study exposes the prevalent issue of factual errors in chart captions generated by LVLMs, such as GPT-4V, and introduces Chocolate to scrutinize these errors. We establish the Chart Caption Factual Error Correction task to propel the creation of trustworthy captioning systems and present ChartVE, an evaluation model surpassing LVLMs in mirroring human assessments of caption factuality. Our two-stage correction framework, C2TFec, provides an interpretable means of improving caption factuality by transforming visual data into structured tables for more faithful error corrections. Our work marks an essential step in ensuring verifiable and trustworthy chart captions. Future directions include extending our approach to multimodal contexts beyond charts, developing more sophisticated error detection and correction algorithms, and creating datasets covering a broader range of visual content.
@@ 9 Limitations
In addition, most charts in the datasets we used are line plots and bar plots. Future efforts can extend our work with additional analyses for other types of charts, such as scatter plots, violin plots, and distribution plots.
@@ Appendix A Further Discussions
Main Results
We see that C2TFec outperforms the pipeline approaches of DePlot/$\text{DePlot}_{\text{CFT}}$ + GPT-4 across the board. While both methods utilize an intermediate tabular representation and leverage GPT-4 for language generation/correction, C2TFec employs a superior chart-to-table conversion model with much more comprehensive training datasets. This results in extracted tables that more faithfully capture the underlying chart data, better facilitating the downstream factual error correction. C2TFec also requires a relatively small number of edits to captions according to Levenshtein distance, making focused changes to improve factuality while minimizing revisions. An example output from C2TFec is shown in Figure 4. By comparison, the proprietary LVLM Bard produces corrected captions requiring 127.83 as many character-level edits on average. This signals excessive rewriting rather than targeted error correction.

### v2

@@ 4 Methodology
In correcting factual errors in generated captions, we propose C2TFec, a two-step, interpretable framework, as shown in Figure 2. C2TFec first transforms input charts into data tables (Section 4.1), then rectifies errors in the caption using the tabular data (4.2). This framework is motivated by our analysis on “DePlot + GPT-4”, which shows that a notable proportion of errors in caption generation originated from the DePlot component. To mitigate this, we develop a stronger chart-to-table model based on UniChart, significantly improved with expansive fine-tuning datasets. The advantage of C2TFec is its ability to harness the reasoning strengths of GPT-4 to faithfully correct errors, boosting caption factuality. [^6] Here, we do not consider approaches based on LVLMs due to their tendency towards factual errors.
@@
Table 4: Correction performance of different models on the Chocolate dataset. ChartVE measures factuality by computing the entailment probability from each chart to the corresponding caption sentences. GPT-4V and Bard, when used as evaluation metrics, rate each chart caption as factually consistent with the chart or not. Levenshtein computes the edit distance between the corrected caption and the original caption (denoted as “N/A”). Metric scores are shown separately for each of the three data splits based on captioning model source. The highest and second highest performing models per evaluation metric and split are highlighted in boldface and underlines respectively.
@@ 4.1 Chart-To-Table Conversion
The training data for our chart-to-table model is sourced from datasets including VisText, Chart-to-Text, ChartQA, and PlotQA, where we repurpose original charts and underlying data tables for our model’s training. We collected a total of 65K instances with a train:dev:test split of 61:2:2. Similar to DePlot Liu et al. (2023a), our model is also trained to generate chart titles, enhancing its ability to contextualize the data represented in table form. Let $\mathcal{M}$ denote our proposed model. For a given chart figure $\mathcal{E}$, the model autoregressively generates a chart title $\mathcal{T}$ and a corresponding table $\mathcal{A}$ (i.e. $\mathcal{T},\mathcal{A}=\mathcal{M}(\mathcal{E})$).
@@ 4.2 Table-based Error Rectification
With the input chart now converted into structured tabular data, the second phase uses the reasoning capacity of LLMs to address the factual inconsistency between $\mathcal{C}$ and the generated table $\mathcal{A}$. Here, we use GPT-4 as the LLM. GPT-4 first provides an explanatory breakdown of detected factual errors in $\mathcal{C}$ based on the table contents. It then uses this explanation to produce a corrected caption $\mathcal{\hat{C}}$. This transparent process enables users to validate the reasoning behind each correction.
C2TFec separates the factual verification from language generation, taking advantage of the complementary strengths of separate vision and language models tailored to their respective domains. The symbolic table representation acts as a bridge to enhance and validate factual consistency in chart captions.
@@ 5 Experimental Settings
To assess C2TFecs ability in factual error correction for chart captions, we experiment on the Chocolate dataset.
@@
Our Chocolate dataset includes 1,187 chart-caption pairs with factually consistent and inconsistent captions, as detailed in Section 2. It is split into Lvlm, Llm, and Ft, reflecting the diversity of models that generated the captions.
@@
Since Chocolate does not comprise training data, we compare C2TFec against zero-shot capable LVLMs and LLMs, including LVLMs, LLaVA-1.5-13B, GPT-4V, Bard, as well as DePlot + GPT-4. For a fairer comparison between our approach and DePlot, we continue fine-tuning DePlot for an additional 5,000 steps on VisText, an approach which has been shown effective for adapting models to unseen domains Huang et al. (2023b). We denote this model as $\text{DePlot}_{\text{CFT}}$. The prompts used for each model are described in Appendix E.
@@
We assess the factual consistency between corrected captions and input charts using ChartVE, GPT-4V, and Bard, according to our recommendations in Section 3.2. In addition, since corrections should be made with as few edits as possible, we measure the number of edits using the Levenshtein distance Levenshtein et al. (1966).
@@ 6.1 Main Results
The results in Table 4 demonstrate that our C2TFec achieves the best performance for factual consistency on the Lvlm and Ft splits, and takes the second place on the Llm split. This indicates that the two-step process of first transforming charts into structured data tables and then rectifying factual inconsistencies using table-caption alignment is an effective strategy.
Additionally, we see that C2TFec outperforms the pipeline approaches of DePlot/$\text{DePlot}_{\text{CFT}}$ + GPT-4 across the board. While both methods utilize an intermediate tabular representation and leverage GPT-4 for language generation/correction, C2TFec employs a superior chart-to-table conversion model with much more comprehensive training datasets. This results in extracted tables that more faithfully capture the underlying chart data, better facilitating the downstream factual error correction. C2TFec also requires a relatively small number of edits to captions according to Levenshtein distance, making focused changes to improve factuality while minimizing revisions. An example output from C2TFec is shown in Figure 4. By comparison, the proprietary LVLM Bard produces corrected captions requiring 127.83 as many character-level edits on average. This signals excessive rewriting rather than targeted error correction. After manually inspecting Bard’s outputs, we found the reason is that Bard oftentimes try to improve the fluency of the caption by paraphrasing. Hence, it makes more edits to the generated captions.
Bard’s underperformance on the Lvlm split and its negative correlation with human judgments of factuality, as shown in Table 3, implies its unreliability in detecting errors in chart captions. Additionally, when used as an evaluator, GPT-4V tends to assign high factuality scores to its own corrected outputs on all three splits (see Table 12), while other metrics show GPT-4V lagging behind C2TFec. This suggests GPT-4V may suffer from the self-enhancement bias Zheng et al. (2023), overestimating its own performance when used for evaluation. We thus perform human evaluations in Section 6.2 to verify the effectiveness of our approach.
@@ 6.2 Human Evaluation
Our human assessments focus on comparing C2TFec with GPT-4V by using the same annotation tasks detailed in Section 2 for factual error identification, with the same annotators evaluating. We sampled 30 charts from each split of Lvlm, Llm, and Ft. For each chart, human judges are presented with a caption generated by one of the models.
Figure 3 demonstrates C2TFecs superiority in multiple error categories, especially with a substantial decrease in Value Errors, over 20% better in the Lvlm and Llm splits, and halving the overall error rate compared to GPT-4V. C2TFec virtually eliminated Trend Errors, highlighting its strong error correction ability, particularly for axes-related errors like Label, Value, and Trend errors. A representative comparison is shown in Figure 4. GPT-4V’s shortcomings seem to stem from its failure to accurately infer data point values from charts as evidenced in Figure 7.
In contrast, GPT-4V is better in addressing Out-of-context Errors, involving information out of the chart’s scope. However, GPT-4V seemed challenged in rectifying errors within captions generated by itself, particularly within the Lvlm split. This observation echoes recent findings on LLMs’ inability to self-correct Huang et al. (2023a); Valmeekam et al. (2023), we find that LVLMs also cannot perform self-correction. More importantly, our human evaluation results, combined with our findings in Table 4 and Table 12, reflect that GPT-4V is subject to serious self-enhancement bias. Consequently, although GPT-4V’s capabilities are formidable, we recommend not using them to assess their own outputs.
@@ 7.1 Chart Captioning
Chart captioning is essential for accurately interpreting and communicating the information conveyed by chart images, particularly in news articles and social media, where factuality is imperative to prevent misinformation. While current datasets like FigureQA Kahou et al. (2017), DVQA Kafle et al. (2018), PlotQA Methani et al. (2020), VisText Tang et al. (2023a), and Chart-to-Text Kantharaj et al. (2022) offer chart image descriptions and question-answer pairs to train models, advancements in vision-language models like ChartT5 Zhou et al. (2023), MatCha Liu et al. (2023b), and UniChart Masry et al. (2023) have largely prioritized relevance and fluency over factual accuracy. Our work provides a rigorous characterization of factual errors in chart captioning and comparisons of methods to address this gap. By focusing on faithfulness and correction, we complement the emphasis of prior work and aim to produce more trustworthy chart captions.
@@ 7.2 Factual Error Correction
Prior research in factual error correction has mainly targeted text summarization and fact-checking. Within summarization, the bulk of work has been carried out in the news domain and often involves methods that substitute inconsistent entities from the source text. Some studies have enhanced this approach through entity-replacement reranking techniques Chen et al. (2021), autoregressive models for rewriting and perturbation filtering Cao et al. (2020); Zhu et al. (2021); Adams et al. (2022), and editing strategies that focus on selective deletion Wan and Bansal (2022). In contrast, Fabbri et al. (2022a) employed sentence compression datasets to train their models. More recently, Gao et al. (2023) have expanded the focus of these studies to include dialogue summarization.
Moving to the domain of fact-checking, this area has experienced a flurry of activity, particularly with the increased attention on combating misinformation Fung et al. (2021); Wu et al. (2022); Fung et al. (2022); Huang et al. (2023d, c); Qiu et al. (2023b); Huang et al. (2024); Qiu et al. (2024). Early approaches train a distantly supervised model that involves a masker and a corrector Shah et al. (2020); Thorne and Vlachos (2021). Thorne and Vlachos (2021) made significant strides by developing the first factual error correction dataset for fact-checking, thus enabling fully supervised training for error correctors. Recently, Huang et al. (2023b) propose an interpretable framework that breaks down the process of factual error correction into individual components. Our study builds on these insights and extends them to a multimodal context, which challenges models to understand the chart images and the consistency between different modalities.
@@ 8 Conclusion
Our study exposes the prevalent issue of factual errors in chart captions generated by various chart captioning models and introduces Chocolate to scrutinize these errors. We establish the Chart Caption Factual Error Correction task to propel the creation of trustworthy captioning systems and present ChartVE, an evaluation model surpassing LVLMs in mirroring human assessments of caption factuality. Our two-stage correction framework, C2TFec, provides an interpretable means of improving caption factuality by transforming visual data into structured tables for more faithful error corrections. Our work marks an essential step in ensuring verifiable and trustworthy chart captions. Future directions include extending our approach to multimodal contexts beyond charts, developing more sophisticated error detection and correction algorithms, and creating datasets covering a broader range of visual content.
9 Ethical Considerations
Text generation models pre-trained on information from the Web are known to demonstrate various biases. Despite the primary focus on models and datasets that represent the English-speaking population’s culture, manual examinations of the Chocolate dataset reveal no evidence of biases related to gender, age, race, or other socioeconomic factors Qiu et al. (2023a); Wang et al. (2024).
In Section 2 and Section 6.2, we recruited annotators to assess the factual consistency of chart captions. The annotators were fairly compensated for their efforts, as detailed in Appendix B. During the annotation process, we made provisions for open communication, allowing the annotators the flexibility to work at their preferred pace and the freedom to withdraw from the project at any point. Additionally, we took measures to protect the anonymity of the contributors by excluding any personally identifiable information from the dataset.
@@ 10 Limitations
In addition, charts in the datasets we used are mostly line plots and bar plots. Future efforts can extend our work with additional analyses for other types of charts, such as violin plots and distribution plots.
Moreover, our investigation centered on the factuality of machine-generated chart captions, particularly those produced by LVLMs. We chose not to examine captions written by humans, as our primary objective was to highlight concerns regarding the reliability of automated chart captioning, a tool on which there is an increasing dependence for humans. The analysis of human-generated chart captions represents another avenue for future research, which could offer valuable comparisons and insights into the effectiveness of automated versus human captioning practices.
Acknowledgement
This research is based upon work supported by U.S. DARPA SemaFor Program No. HR001120C0123, DARPA ECOLE Program No. HR00112390060, and the Molecule Maker Lab Institute: an AI research institute program supported by NSF under award No. 2019897 and No. 2034562. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein. Hou Pong Chan was supported in part by the Science and Technology Development Fund, Macau SAR (Grant Nos. FDCT/060/2022/AFJ, FDCT/0070/2022/AMJ) and the Multi-year Research Grant from the University of Macau (Grant No. MYRG2020-00054-FST).
@@ Appendix A Further Discussions
Error Typology
Our error typology was derived from our systematic preliminary analysis of around 50 generated captions. This exploratory phase allowed us to identify common error patterns and categories emergent from the data itself. Moreover, our methodology was enriched by drawing on insights from prior works such as Chart-to-Text, ChartT5, and VisText.
Model Usage
The GPT-4V version we used was gpt-4-vision-preview. We query GPT-4V via API and the cost for using GPT-4V to produce captions for our CHOCOLATE dataset is less than \$50 USD. For the Bard model, we obtain its outputs during October 2023. There is no cost using Bard since we access it via its web interface. Both models are prompted in single conversation, as seen in Figure 10.
@@
[…]
Table 5: Sentence-level error counts and error rates.
[…]
Table 6: Caption-level error counts and error rates.
Furthermore, in Figure 1, the error rates are computed as the number of such errors divided by the number of sentences. While this provides an overview of the frequency for each error, it does not indicate the likelihood of a value/label/trend/magnitude-related mention in the generated captions being factual. This limitation can result in an underrepresentation of certain error types – for instance, the infrequent occurrence of Magnitude Errors as shown in Figure 1 is more a consequence of the scarcity of magnitude-related mentions in the captions rather than an indication of the models’ superior trend variance comprehension. To address this, we sample 30 generated captions for each model from each dataset and compute another error rate as the number of sentences containing such non-factual mentions over the number of sentences containing such mentions. The results are shown in Table 7. The outcomes corroborate the observations in Section 2.2, while Table 7 offers a supplementary perspective on model performance.
Meta-evaluation Results
For the text-based metrics presented in Table 3, they both perform weakly in determining the factuality of the generated caption. This is largely because charts often contain much denser information compared to the corresponding reference. As a result, text-only factuality metrics are unsuitable for assessing factual consistency between charts and captions. Additionally, the negative values in Table 3 signal instances where the performance of the methods inversely correlates with human judgments, indicating their poor effectiveness in serving as evaluation metrics for identifying factual errors in generated captions. Furthermore, we experimented with ROC AUC as an additional meta-evaluation metric. As shown in Table 9, our original conclusion with Kendall’s Tau still holds.
Understanding The Upper Bound
We seek to understand the performance upper bound of our proposed two-stage framework by replacing generated tables with ground-truth data tables. Since the ground-truth data tables in Pew are not available, we experiment with only the instances from the VisText dataset. The results are demonstrated in Table 8.
Ablation Studies
We compare our chart-to-table component against DePlot, $\text{DePlot}_{\text{CFT}}$, and GPT-4V. Since the Pew split of Chart-to-Text does not contain ground truth labels, we show performance on the VisText test set. We use RMS-F1 proposed in Liu et al. (2023a) as the evaluation metric. From Table 11, we see that our Chart-to-Table component significantly outperforms all baselines, indicating its effectiveness in converting charts into their underlying data tables.
Figure 5: Human annotation interface for our data collection discussed in Section 2. Examples of each type of error from Table 1 are also displayed in the annotation interface. We were not able to show these examples in this figure due to space limits.